OpenStudy (anonymous):

Use Cramer’s rule to solve for y in the following system of equations. x + 4y – 2z = 3 x + 3y + 7z = 1 2x + 9y + z = 8

OpenStudy (anonymous):

x + 4y – 2z = 3 , time(-1): -x-4y+2z=-3 -x- 4y+2z =-3 x + 3y + 7z = 1 ADD--------------- -y +9z = -2 x + 3y + 7z = 1, time (-2): -2x-6y-14z =-2 -2x-6y-14z =-2 2x + 9y + z = 8 ADD------------------ 3y -13z =6 -y +9z = -2, time(3): -3y +27z =-6 -3y +27z =-6 3y -13z =6 ADD---------------- 14z=0 z=0 Substitution z =0 find y 3y -13z =6 3y -13(0) =6 3y= 6 y =6/ 3 y=2 Substitution z =0 , y =2 find x x + 4y – 2z = 3 x = - 4y + 2z + 3 x = - 4(2) + 2(0) + 3 x = -8 +3 x = -5

OpenStudy (anonymous):

\[\text{LinearSolve}\left[\left( \begin{array}{ccc} 1 & 4 & -2 \\ 1 & 3 & 7 \\ 2 & 9 & 1 \end{array} \right),\left( \begin{array}{c} 3 \\ 1 \\ 8 \end{array} \right)\right]=\{-5,2,0\} \]

OpenStudy (anonymous):

http://www.youtube.com/watch?v=taBHTo8sviM